# Sufficient conditions for QMC analysis of finite elements for

### v1

@@ 4.3 Continuous finite elements
Continuous (also classic, conforming, or standard) finite elements approximate the solution by $u_{h}\in U_{h}\subset H^{1}(D)$ with
@@
Setting $\boldsymbol{q}_{h}=-a^{-1}(\boldsymbol{y})\nabla u_{h}(\boldsymbol{y})\in Q_{h}=L^{2}(\Omega)$ in the interior of all $E\in{\mathcal{E}_{h}}$ and $\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}=\|\sqrt{a(\boldsymbol{y})}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{L^{2}(D)}$ allows us to employ our general framework—note that in this case $Q_{h}$ is not finite dimensional.
@@
$\|\boldsymbol{q}_{h}(\boldsymbol{y})\|^{2}_{Q_{h}}=\int_{D}fu_{h}(\boldsymbol{y})\;\textup{d}\boldsymbol{x}\leq C_{S}\sqrt{\overline{a}(\boldsymbol{y})}\|f\|_{L^{2}(D)}\|\boldsymbol{q}_{h}(\boldsymbol{y})\|^{2}_{Q_{h}}.$
@@
$\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|^{2}_{Q_{h}}$ $=-\sum_{\boldsymbol{0}\neq\boldsymbol{m}\leq\boldsymbol{\nu}}\begin{pmatrix}\boldsymbol{\nu}\\ \boldsymbol{m}\end{pmatrix}([\partial^{\boldsymbol{m}}_{\boldsymbol{y}}a(\boldsymbol{y})]\partial^{\boldsymbol{\nu}-\boldsymbol{m}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y}),\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y}))_{D}$
$\leq\sum_{\boldsymbol{0}\neq\boldsymbol{m}\leq\boldsymbol{\nu}}\begin{pmatrix}\boldsymbol{\nu}\\ \boldsymbol{m}\end{pmatrix}\left\|\frac{\partial^{\boldsymbol{m}}_{\boldsymbol{y}}a(\boldsymbol{y})}{a(\boldsymbol{y})}\right\|_{L^{\infty}(D)}\|\partial^{\boldsymbol{\nu}-\boldsymbol{m}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}.$
@@ 5 Numerical experiments
(cf., e.g., [37, p. 16]) and the convention that $\operatorname{sign}(0)=0$.
We use the off-the-shelf generating vector [28, lattice-32001-1024-1048576.3600] with a total amount of $2^{m}$, $m=2,\dots,14$ cubature points with $R=16$ random shifts. Although this generating vector has not been obtained using the CBC algorithm with the weights derived in [17] and Theorem 3.12, we found this off-the-shelf generating vector to perform well, yielding the optimal rate of convergence. Thus, this confirms our analytical findings without excluding that there might be even better lattice rules. In practice, the off-the-shelf lattice rule performs similarly to tailored lattice rules. The finite element discretization uses the lowest order Raviart–Thomas discretization on an unstructured simplicial mesh. The file that needs to be run to reproduce our results can be found in [27]; it uses NGSolve [38].
@@
For the lognormal case (Figure 1, (C)), we observe that the root mean square error for the primal unknown $u_{h}(\boldsymbol{y})$ decreases with a rate of $0.90$ and the root mean square error for the gradient $a(\boldsymbol{y})\boldsymbol{q}_{h}(\boldsymbol{y})$ decreases with a rate of $0.90$. Thus, these two quantities seem to decrease optimally. However, the convergence rate for the flux unknown $\boldsymbol{q}_{h}(\boldsymbol{y})$ is slightly better and lies at $0.91$. We also see that the root mean square errors for the primal unknown and the flux are similar. In contrast, the respective error for the gradient is larger and highlights that considering only $2^{i}$ with $i\geq 8$ points yields convergence rates of 0.94 for $u_{h}(\boldsymbol{y})$, 0.96 for $\nabla u_{h}(\boldsymbol{y})$, and 0.98 for $\boldsymbol{q}_{h}(\boldsymbol{y})$ indicating that our QMC error bounds are accurate for large sample sizes. The lognormal Gevrey-1.25 case (Figure 1, (D)) converges with orders 0.98 for gradient and primal unknown and with order 0.96 for the flux unknown.
@@ 6 Conclusions
A limitation of our framework is numerical methods that need a rich enough stabilization to approximate the quantity of interest stably, such as (interior penalty) discontinuous Galerkin or discontinuous Petrov–Galerkin methods. Such methods typically violate (A3) and need individualized analysis techniques; see [26] for discontinuous Galerkin.
@@
    More general equations. A careful revision of the arguments in Sections 2 and 3 reveals that (1) is never used in the QMC analysis. Thus, we could conduct the same analysis if assumptions (A2)(A5) hold for any other (partial differential) equation.
@@
    where we used the fact that $\ell!(|\mathfrak{u}|-\ell)!\leq|\mathfrak{u}|!$ and especially $(\ell!(|\mathfrak{u}|-\ell)!)^{\sigma}\leq(|\mathfrak{u}|!)^{\sigma}$ for $\sigma\geq 1$. This estimate can be used to shorten and improve the estimates of $t_{\mathfrak{u}}$ in the proofs of Theorems 3.8 and 3.12: Repeating the subsequent arguments in the respective proofs allows to obtain the convergence results with $r=1$ instead of $r=2$ for quadratic quantities of interest.

### v2

@@ 4.3 Continuous finite elements
Continuous (also classic, conforming, or standard) finite elements approximate the solution by $u_{h}\in U_{h}$ (where $U_{h}\subset H^{1}(D)$ is an element-wise polynomial finite element space) with
@@
Setting $\boldsymbol{q}_{h}=-a^{-1}(\boldsymbol{y})\nabla u_{h}(\boldsymbol{y})\in\mathcal{W}_{h}=L^{2}(\Omega)$ in the interior of all $E\in{\mathcal{E}_{h}}$ and $\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}=\|\sqrt{a(\boldsymbol{y})}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{L^{2}(D)}$ allows us to employ our general framework—note that in this case $\mathcal{W}_{h}$ is not finite dimensional.
@@
$\|\boldsymbol{q}_{h}(\boldsymbol{y})\|^{2}_{\mathcal{W}_{h}}=\int_{D}fu_{h}(\boldsymbol{y})\;\textup{d}\boldsymbol{x}\leq C_{S}\sqrt{\overline{a}(\boldsymbol{y})}\|f\|_{L^{2}(D)}\|\boldsymbol{q}_{h}(\boldsymbol{y})\|^{2}_{\mathcal{W}_{h}}.$
@@
$\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|^{2}_{\mathcal{W}_{h}}$ $=-\sum_{\boldsymbol{0}\neq\boldsymbol{m}\leq\boldsymbol{\nu}}\begin{pmatrix}\boldsymbol{\nu}\\ \boldsymbol{m}\end{pmatrix}([\partial^{\boldsymbol{m}}_{\boldsymbol{y}}a(\boldsymbol{y})]\partial^{\boldsymbol{\nu}-\boldsymbol{m}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y}),\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y}))_{D}$
$\leq\sum_{\boldsymbol{0}\neq\boldsymbol{m}\leq\boldsymbol{\nu}}\begin{pmatrix}\boldsymbol{\nu}\\ \boldsymbol{m}\end{pmatrix}\left\|\frac{\partial^{\boldsymbol{m}}_{\boldsymbol{y}}a(\boldsymbol{y})}{a(\boldsymbol{y})}\right\|_{L^{\infty}(D)}\|\partial^{\boldsymbol{\nu}-\boldsymbol{m}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}.$
@@ 5 Numerical experiments
(cf., e.g., [qi1996general, p. 16]) and the convention that $\operatorname{sign}(0)=0$.
We use the off-the-shelf generating vector [lattice, lattice-32001-1024-1048576.3600] with a total amount of $2^{m}$, $m=2,\dots,14$ cubature points with $R=16$ random shifts. Although this generating vector has not been obtained using the CBC algorithm with the weights derived in [GrahamKNSSS15] and Theorem 3.14, we found this off-the-shelf generating vector to perform well, yielding the theoretically expected rate of convergence. Thus, this confirms our analytical findings without excluding that there might be even better lattice rules. In practice, the off-the-shelf lattice rule performs similarly to tailored lattice rules. The finite element discretization uses the lowest order Raviart–Thomas discretization on an unstructured simplicial mesh. The file that needs to be run to reproduce our results can be found in [KaarniojaRG25]; it uses NGSolve [Schoberl14].
@@
For the lognormal case (Figure 1, (C)), we observe that the root mean square error for the primal unknown $u_{h}(\boldsymbol{y})$ decreases with a rate of $0.90$ and the root mean square error for the gradient $a(\boldsymbol{y})\boldsymbol{q}_{h}(\boldsymbol{y})$ decreases with a rate of $0.90$. Thus, these two quantities seem to decrease at the theoretically expected rate. However, the convergence rate for the flux unknown $\boldsymbol{q}_{h}(\boldsymbol{y})$ is slightly better and lies at $0.91$. We also see that the root mean square errors for the primal unknown and the flux are similar. In contrast, the respective error for the gradient is larger and highlights that considering only $2^{i}$ with $i\geq 8$ points yields convergence rates of 0.94 for $u_{h}(\boldsymbol{y})$, 0.96 for $\nabla u_{h}(\boldsymbol{y})$, and 0.98 for $\boldsymbol{q}_{h}(\boldsymbol{y})$ indicating that our QMC error bounds are accurate for large sample sizes. The lognormal Gevrey-1.25 case (Figure 1, (D)) converges with orders 0.98 for gradient and primal unknown and with order 0.96 for the flux unknown.
@@ 6 Conclusions
A limitation of our framework is numerical methods that need a rich enough stabilization to approximate the quantity of interest stably, such as (interior penalty) discontinuous Galerkin or discontinuous Petrov–Galerkin methods. Such methods typically violate 3 and need individualized analysis techniques; see [KaarniojaR24] for discontinuous Galerkin.
@@
    More general equations. A careful revision of the arguments in Sections 2 and 3 reveals that (1) is never used in the QMC analysis. Thus, we could conduct the same analysis if Assumptions 25 hold for any other (partial differential) equation.
@@
    where we used the fact that $\ell!(|\mathfrak{u}|-\ell)!\leq|\mathfrak{u}|!$ and especially $(\ell!(|\mathfrak{u}|-\ell)!)^{\sigma}\leq(|\mathfrak{u}|!)^{\sigma}$ for $\sigma\geq 1$. This estimate can be used to shorten and improve the estimates of $t_{\mathfrak{u}}$ in the proofs of Theorems 3.10 and 3.14: Repeating the subsequent arguments in the respective proofs allows to obtain the convergence results with $r=1$ instead of $r=2$ for quadratic quantities of interest.